OpenStudy (anonymous):

Use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth. x² + 3 = –4x

OpenStudy (anonymous):

Hint... use the quadratic equation... \[\LARGE x_{1/2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] and substitute a=1 b=4 c=3

OpenStudy (anonymous):

Alright Thank you!

OpenStudy (anonymous):

Glad to help :)